AgNO3+NH4Cl --> AgCl+NH4NO3. If you have 50.26g of AgNO3, how many grams of NH4NO3 can be produced?
1
Expert's answer
2018-01-11T13:27:06-0500
AgNO3+NH4Cl = AgCl+NH4NO3
n(AgNO3) = n(NH4NO3)
n(AgNO3) = 50,26 / 169,87 = 0,296 mol = n(NH4NO3)
m(NH4NO3) = n(NH4NO3) * Mr(NH4NO3)
m(NH4NO3) = 0,296 * 80 = 23,67 g